OpenStudy (anonymous):

How to factorise 6x^2+5x+1

OpenStudy (anonymous):

choose 2 numbers which a*b=6 & a+b=5

OpenStudy (anonymous):

(ax+1)(bx+1)

OpenStudy (anonymous):

Thanks a lot

OpenStudy (perl):

Suppose you wanted to factor: 10x^2 + 29x + 21 First multiply 10*21 = 210 Find two factors that multiply to 210 and add up to 29 (the middle number). two factors are 15,14 now decompose the middle term 29x 10x^2 + 15x + 14x + 21 now factor by grouping 5x(2x + 3) + 7(2x+3) (2x+3)(5x+7)

OpenStudy (anonymous):

yourwelcome